The grade point averages (GPA) of 18 randomly selected college students are used to estimate the mean GPA of the college students. The GPAs from the sample are as follows:

2.3       3.3       2.6       1.8       0.2 3.1       4.0       0.7          3.1

2.3       2.0       3.1       3.4       1.3       2.6       2.6       3.7          2.2

  1. (5 points) Is it justified to use of the standard normal distribution to construct the confidence interval? Explain.
  2. (10 points) If the population is assumed to be normally distributed, construct a 98% confidence interval for the population mean GPA. Show your work in detail and use 3 decimal digits.
Answer #1

I have answered the question below

Please up vote for the same and thanks!!!

Do reach out in the comments for any queries

Answer:

x x2
2.3 5.29
3.3 10.89
2.6 6.76
1.8 3.24
0.2 0.04
3.1 9.61
4 16
0.7 0.49
3.1 9.61
2.3 5.29
2 4
3.1 9.61
3.4 11.56
1.3 1.69
2.6 6.76
2.6 6.76
3.7 13.69
2.2 4.84
∑x=44.3 ∑x2=126.13

Mean *x={xn =2.3+3.3+2.6+1.8+0.2+3.1+4+0.7+3.1+2.3+2+3.1+3.4+1.3+2.6+2.6+3.7+2.2/18 =44.3/18 =2.4611 Sample Standard deviatio

Margin of error = 2.567 The 98% confidence interval estimate of the population mean is, T-E < < + E 2.461 - 2.567 < p < 2.461